Use the following information for questions 27 – 29. Bill is a corn farmer in the Texas Panhandle. He has a 10 year average corn production of 25,000 bushels on his farm. At no time in the past 5 years has that production dropped below 15,000 bushels. On March 5, Bill notices the December CBOT corn futures are trading at $4.178 per bushel. This is a much higher price than Bill has seen in the past and he wants to lock in his profits on his minimum production level of 15,000 bushels. Margin requirements are $1,500 per contract and a standard corn contract covers 5,000 bushels. (Explain)

27. Which of the following best describes Bill’s position on March 5 and his hedging strategy?

a. Since Bill is short in the underlying asset, he should buy 5 futures contracts. b. Since Bill is short in the underlying asset, he should buy 3 futures contracts. c. Since Bill is long in the underlying asset, he should buy 3 futures contracts. d. Since Bill is long in the underlying asset, he should short 3 futures contracts. e. Since Bill is long in the underlying asset, he should short 5 futures contracts.

28. How much margin must Bill deposit or receive when opening the future’s hedge?

a. Bill must deposit $4,500 since he is short 3 contracts. b. Bill will receive $4,500 since he is long 3 contracts. c. Bill must deposit $7,500 since he is short 5 contracts. d. Bill will receive $7,500 since he is long 5 contracts. e. Bill must deposit $62,670 since this is the initial value of his hedged position.

29. On October 10, Bill sells his entire 28,000 bushels of harvested and dried corn at the spot price of $4.785 per bushel and at the same time he closed his December futures contracts at $5.694 per bushel. How much profit or loss did Bill make on the future’s contracts?

a. $22,740 loss b. $13,635 loss c. $9,105 profit d. $13,635 profit e. $22,740 profit
Answer #1

27) We see that Bill is a producer of Corn Bushel, which is the underlying asset. One future contract is of 5000 Bushel. Therefore we see that he needs to sell three contracts (15000/5000 =3 Contracts)

So Bill being a producer of Bushel is long on the underlying asset, which means he already possess the underlying asset and he therefore needs to short on 3 Futures Contract, which means option no. d.

28) Since in the above answer we computed that Bill is short on 3 contract and therefore he needs to submit the margin money on three contract.(1500*3=$4500). Hence the correct answer is a

29) Bill had contracted to sell the bushel @$4.178 per bushel, now on expiry of the contract the spot rate is $5.694 Per bushel. Therefore bills loss = (4.178-5.694)*15000 Bushel = $-22740 Loss. Hence the correct answer is a